### v1

@@ 5.1 Stacking processes
For each filament, we first merge the event files of TM8 and shift the Pulse Invariant of each event by a factor of $1+z$. Then, we use srctool to extract the source and local background spectra and the vignetting corrected source ancillary response file (ARF). The source and local background extraction regions are defined as 0–10 Mpc and 10–20 Mpc transverse distance to the filament spine, and whose region masks for spectral extraction are created by constructing images using and excluding sources in Table 1. For each ARF, we first correct it by multiplying the foreground absorption curve given the averaged Galactic $n_{\mathrm{H}}$ value (HI4PI Collaboration et al. 2016) in the extraction region and the TBABS absorption model (Wilms et al. 2000), and then shift it by a factor of $1+z$ to the source rest frame. The response matrix file (RMF) is taken from the calibration database and then shifted to the source rest frame by adopting a two-dimensional interpolation strategy on both axes of induced photon energy and resulting spectral channels (see Bulbul et al. 2014, for further details).
We use tools addspec, addarf, and addrmf in FTOOLS [^3] http://heasarc.gsfc.nasa.gov/ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014) Heasarc) to stack the spectra, ARFs and RMFs, respectively. We note that some keywords in the headers of srctool products have different definitions from those in other X-ray satellite data products [^4] See https://erosita.mpe.mpg.de/edr/DataAnalysis/srctool_doc.html#Output_files for the details of the srctool products.. In short, the EXPOSURE value of a spectrum $t_{\mathrm{EXP}}$ is the total time of an extraction region on a detector, instead of the averaged exposure time in the region $t_{\mathrm{ave}}$; BACKSCAL value $\Omega_{\mathrm{BS}}$ is the averaged intersection of the extraction region and the detector area; the geometric solid angle of the extraction region $\Omega_{\mathrm{REG}}$ has a keyword REGAREA, and $\Omega_{\mathrm{REG}}/\Omega_{\mathrm{BS}}=t_{\mathrm{EXP}}/t_{\mathrm{ave}}$. Furthermore, the normalization of an ARF curve is scaled by a factor of $t_{\mathrm{ave}}/t_{\mathrm{EXP}}$. Therefore, we add up all individual source spectra to obtain a stacked source spectrum.
$N_{\mathrm{stacked,src}}(E)=\sum_{i}N_{\mathrm{src},i}(E),$ (9)
and scale each local background spectrum to match the extraction solid angle of the corresponding source spectrum to get the stacked background spectrum
@@
The blueshifted response files at the rest frame corresponding to each filament are then later stacked using FTOOLS addarf and addrmf to obtain one final stacked response file. These files are used in the modeling of the stacked spectrum described in the section below.
One clear advantage of the spectral stacking method in the rest frame is the reduction of the effect on the instrumental and background features on the final net background-subtracted source spectrum as described in (Bulbul et al. 2014). The energies of the instrumental and celestial X-ray foreground/background features, which are frozen in the detector frame, are smoothed over the redshift range covered by the sample, therefore reducing their overall effect on the net spectrum. The spectral stacking method is particularly well suited for obtaining a spectrum of faint sources, such as cosmic filaments.
@@ 5.2 Stacked spectrum and spectral fitting
We stacked the 0–10 Mpc transverse distance region spectra as the source spectrum and the 10–20 Mpc transverse distance region spectra as the local background region spectrum. The stacked background spectrum is smoothed using a Wiener filter with a kernel size of 5 channels for direct subtraction. Additionally, we compare the 6–9 keV count rate of the two spectra and store the ratio as the backscal in the header for further correction. The stacked net source spectrum is plotted in Fig. 9.
@@
We use XSPEC version 12.13.0 and the package BXA [^5] https://github.com/JohannesBuchner/BXA (Buchner et al. 2014) for spectral analysis. The spectrum is first optimally binned (Kaastra & Bleeker 2016) using the tool ftgrouppha and then loaded into XSPEC. We use the 0.5–6.0 keV range for analysis and adopt C-stat (Kaastra 2017) to derive the fitting likelihood. The BXA package uses the nested sampling algorithm UltraNest (Buchner 2016, 2019, 2021) for Bayesian inference.
@@
We fit the spectrum using four components that account for the detected WHIM, unmasked halo gas, unmasked AGN, and unmasked XRB. The spectral model of the three unresolved sources has been described in Sect. 4.1. We calculate the expected normalization ratios among the three components based on the modeled unresolved source fractions in Sect. 4.2 and coupled the ratios during fitting. To model the WHIM component, we use a single temperature (1-T) APEC model with fixed abundance $Z=0.2Z_{\sun}$. This value is based on the reported $\lesssim 0.3Z_{\odot}$ abundance in cluster outskirts by both observations and numerical simulations (Bulbul et al. 2016; Mernier et al. 2017; Biffi et al. 2017) and $>0.1Z_{\odot}$ in $z<1$ filaments from numerical simulations (Martizzi et al. 2019). This model combination can well fit the $E>0.7$ keV spectrum range with a best-fit parameter $kT=0.58\pm 0.10$ keV and a fitting statistics $c_{\mathrm{stat}}/\mathrm{d.o.f.}=105/111$. We also tried to fix the metalicity to $0.1Z_{\sun}$ and $0.3Z_{\sun}$ to test the impact of metalicity assumption and found there is has no impact on the best-fit temperature.
@@
Recently, Vijayan & Li (2022) demonstrates that fitting a multi-temperature nature spectrum using a 1-T model will lead to an overestimation of the temperature. It explains that our best-fit temperature using a 1-T model does not agree with the majority of the gas phases around $10^{6}$ K predicted by simulations (e.g., Martizzi et al. 2019; Galárraga-Espinosa et al. 2021). In our stacked spectrum, we observe an excess feature in 0.5–0.6 keV, which agrees with the energy range of $\sim 0.56$ keV O VII lines emitted by cooler phases of gas around $10^{6.3}$ K ($\sim 0.17$ keV), suggesting the presence of cooler phases. The gas in these temperature phases only contributes to lines below 0.7 keV, while the fitting using a 1-T model tends to maximize the likelihood around the peak of the observed spectrum, which is in the 0.8–0.9 keV energy range, dominated by several Fe XVII lines emitted by $10^{6.8}$ K gas.
@@
We also try to fit the spectrum using a multi-temperature model. We adopt a log-normal temperature distribution DEM model with a peak at $\log(T/\mathrm{K})=5.9$ and $1\sigma$ width of 0.45 dex, which is suggested by our analysis of numerical simulations (see Sect. 6.2 later). We use a simplified assumption of CIE condition and $0.2Z_{\sun}$ abundance of the WHIM. The log-normal DEM model can better fit the stacked spectrum in the 0.5–0.6 keV energy range (see the bottom panel of Fig. 9), though this fitting shows the feature of a likely deficit around 0.7 keV. It could be due to that the true temperature distribution is more complex than the log-normal distribution we use, which creates discrepancy in some energy ranges.
Figure 9: Left: Rest frame stacked source spectrum of the 0–10 Mpc region from the filament spine in the transverse direction after the subtraction of the local background extracted in the 10–20 Mpc region surrounding the filament. The spectrum is binned to a minimum signal-to-noise ratio of 2.5 for illustration. The combination of a single temperature CIE (red dashed line), and the unmasked sources models (green, blue, and grey) can well fit most of the spectrum, except the significant excess at $E<0.6$ keV. The best-fit temperature of the WHIM gas is $0.58\pm 0.10$ keV. Right: Same as the left panel, but fitted by a log-normal temperature distribution with $\log(\mu/\mathrm{K})=5.9$ and $\sigma=0.45$ dex. The 0.5–0.6 keV energy range can be well fitted by this model, though there is a likely deficit at $\sim 0.7$ keV.
6 The detected phases of WHIM
@@ 6.1 Surface brightness profile modeling
Table 3: Parameters and the corresponding priors for constraining the density profile of the observed phase WHIM.
@@
With the best-fit temperature obtained in the previous section, we fit the density profiles of the warm gas in cosmic filaments. In this section, we use $r$ to denote the physical radius of gas cells/particles to the filament spine in the 3-D space. Based on results from numerical simulations (Cui et al. 2019; Galárraga-Espinosa et al. 2021; Tuominen et al. 2021), the gas in filaments has a multi-phase nature ranging from $10^{3}$ to $10^{7}$ K in temperature and from $10^{-6}$ to $10^{-1}$ cm-3 in hydrogen density. Therefore, we introduce a parameter $f_{\mathrm{vol}}$ to represent the fraction of the volume of the detected gas to the total filament volume. The volume-averaged emissivity of the detected phase gas at a certain radius is
@@
where $n_{\mathrm{H}}$ is the hydrogen density, 1.20 is the $n_{\mathrm{e}}/n_{\mathrm{H}}$ ratio for the plasma at the detected temperature. We use a King-type profile (King 1962), which is widely used to fit central flattened profiles, to model the radial change of both $n_{\mathrm{H}}$ and $f_{\mathrm{vol}}$,
@@
where, $n_{\mathrm{H,0}}$ ($f_{\mathrm{vol,0}}$), $r_{\mathrm{c,n}}$ ($r_{\mathrm{c,fvol}}$), and $\alpha_{\mathrm{n}}$ ($\alpha_{\mathrm{fvol}}$) are the central normalization, core radius, and slope in the outskirts for $n_{\mathrm{H}}$ (or $f_{\mathrm{vol}}$), respectively. Because the radial dependencies of $n_{\mathrm{H}}$ and $f_{\mathrm{vol}}$ have a strong degeneracy, we assume that they have the same core radii, i.e., $r_{\mathrm{c,n}}=r_{\mathrm{c,fvol}}$.
In addition, we make several assumptions. First, we assume a constant baryon density contrast $\Delta_{\mathrm{b,0}}$ at the filament spine for filaments at all redshifts. In this case,
@@
where $\rho_{\mathrm{crit}}$ is the critical density at $z=0$, the factor 1.43 is the ratio between total mass density and hydrogen mass density, $m_{\mathrm{H}}$ is the hydrogen mass, and the redshift evolution comes from the expansion of the physical scale. Second, there is an outer boundary of the filaments $r_{\mathrm{out}}$, beyond which the gas density of our detected phase is zero. In this work, we adopt a value of 10 Mpc. Third, the observed profile is smoothed and broadened due to the position uncertainty of the skeleton constructed from incomplete galaxy samples (e.g., Fig. 11 of Galárraga-Espinosa et al. 2021). Including these assumptions, the projected surface brightness at a given projected radial distance $r_{\mathrm{proj}}$ is
@@
where $l$ is depth along the line of sight; $K_{\mathrm{ph}}$ is the K-correction of photon spectrum given the spectral shape, redshift, and energy band; the term $(1+z)^{-3}$ is the photon dimming factor; $\Lambda$ is the cooling function calculated using the APEC model given the temperature, metallicity and energy band; $A_{\mathrm{CF}}$ is the conversion factor from photon flux to count rate, which is calculated given the on-axis ARF, observation band, observed photon spectrum and Galactic absorption curve.
When computing the fitting statistics, we construct 7817 individual model surface brightness count rate profiles based on the redshift of each filament. We use the solid angle of each filament as the weight to obtain an averaged surface brightness profile as
@@
We use a Bayesian inference approach to obtain posterior distributions of the six parameters, which are $\log\Delta_{\mathrm{b,0}}$, $\log f_{\mathrm{vol,0}}$, $\log r_{\mathrm{c,n}}$, $\alpha_{\mathrm{n}}$, $\alpha_{\mathrm{fvol}}$ and $\log\sigma_{\mathrm{sm}}$, as listed in Table 3. In addition to the fitting statistics $\chi^{2}$, the likelihood function includes an additional constraint from SZ stacking results (Tanimura et al. 2020a), which is necessary to break up the degeneracy between $\Delta_{\mathrm{b,0}}$ and $f_{\mathrm{vol,0}}$. The predicted Compton-$y$ strength is
@@
where $\sigma_{\mathrm{T}}$ is the Thomson scattering cross section, $m_{\mathrm{e}}$ is the electron mass, $c$ is the speed of light. The profiles of Tanimura et al. (2020a) show similar broadening effects as our stacked X-ray profile. Therefore, we also convolved the predicted Compton-$y$ profile with the same Gaussian kernel to create a smoothed Compton-$y$ profile as,
@@
Given that the phases of the WHIM we detect only contribute a fraction of the total stacked SZ signal, we tentatively set the upper limit of the $\log y_{\mathrm{pd,conv}}(0)$ to be -7.5, which is approximately the stacked SZ signal at filament spine reported by (Tanimura et al. 2020a). We use a likelihood term $P(\log y_{\mathrm{pd,conv}}(0))=\mathcal{N}(\mu_{\mathrm{logy}},\sigma_{\mathrm{logy}})$ as the addition constraint. We do not exactly know how much of the total stacked Compton-$y$ is contributed by the X-ray emitting gas, nevertheless, we adopt $\mu_{\mathrm{logy}}=-8.3$ and $\sigma_{\mathrm{logy}}=0.5$ in our analysis. The combined likelihood function is
@@
We use the Markov Chain Monte Carlo (MCMC) method and the python package Zeus [^6] https://zeus-mcmc.readthedocs.io/en/latest/ (Karamanis et al. 2021) to perform Bayesian inference. The marginalized posterior distributions shows that the $\log\Delta_{\mathrm{b,0}}=2.28\pm 0.22$ and $\log f_{\mathrm{vol,0}}=-1.91\pm 0.28$. By using all MCMC chains for calculating the averaged baryon density contrast, we obtain $\log\bar{\Delta}_{\mathrm{b}}=1.88\pm 0.18$, which corresponds to a hydrogen density $4.4^{+2.2}_{-1.5}\times 10^{-5}$ cm-3 at the median redshift of the selected filament sample $z=0.47$.

### v2

@@ 5.1 Stacking processes
For each filament, we first merged the event files of TM8 and shifted the pulse-invariant of each event by a factor of $1+z$. Then, we used srctool to extract the source and local background spectra and the vignetting corrected source ancillary response file (ARF). The source and local background extraction regions are defined as 0–10 Mpc and 10–20 Mpc transverse distance to the filament spine, whose region masks for spectral extraction are created by constructing images using and excluding sources in Table 1. For each ARF, we first corrected it by multiplying the foreground absorption curve given the averaged Galactic $n_{\mathrm{H}}$ value (HI4PI Collaboration et al. 2016) in the extraction region and the TBABS absorption model (Wilms et al. 2000). We then shifted it by a factor of $1+z$ to the source rest frame. The response matrix file (RMF) was taken from the calibration database and then shifted to the source rest frame by adopting a two-dimensional (2D) interpolation strategy on both axes of induced photon energy and resulting spectral channels (see Bulbul et al. 2014, for further details).
We used tools addspec, addarf, and addrmf in FTOOLS [^3] http://heasarc.gsfc.nasa.gov/ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014) Heasarc) to stack the spectra, ARFs, and RMFs, respectively. We note that some keywords in the headers of srctool products have different definitions from those in other X-ray satellite data products [^4] See https://erosita.mpe.mpg.de/edr/DataAnalysis/srctool_doc.html#Output_files for the details of the srctool products.. In short, the EXPOSURE value of a spectrum $t_{\mathrm{EXP}}$ is the total time of an extraction region on a detector, instead of the averaged exposure time in the region $t_{\mathrm{ave}}$; BACKSCAL value $\Omega_{\mathrm{BS}}$ is the averaged intersection of the extraction region and the detector area; the geometric solid angle of the extraction region $\Omega_{\mathrm{REG}}$ has a keyword of REGAREA, along with the ratio $\Omega_{\mathrm{REG}}/\Omega_{\mathrm{BS}}=t_{\mathrm{EXP}}/t_{\mathrm{ave}}$. Furthermore, the normalization of an ARF curve is scaled by a factor of $t_{\mathrm{ave}}/t_{\mathrm{EXP}}$. Therefore, we need to add up all individual source spectra to obtain a stacked source spectrum, as follows:
$N_{\mathrm{stacked,src}}(E)=\sum_{i}N_{\mathrm{src},i}(E).$ (9)
We then scale each local background spectrum to match the extraction solid angle of the corresponding source spectrum to get the stacked background spectrum,
@@
The blueshifted response files at the rest frame corresponding to each filament are later stacked using FTOOLS addarf and addrmf to obtain one final stacked response file. These files are used to model the stacked spectrum described in the section below.
One clear advantage of the spectral stacking method in the rest frame is reducing the effect on the instrumental and background features on the final net background-subtracted source spectrum as described in (Bulbul et al. 2014). The energies of the instrumental and celestial X-ray foreground and background features, frozen in the detector frame, are smoothed over the redshift range covered by the sample, reducing their overall effect on the net spectrum. The spectral stacking method is well suited for obtaining a spectrum of faint sources, such as cosmic filaments.
Refer to captionRefer to caption
Figure 9: Rest frame stacked source spectrum of the 0–10 Mpc region from the filament spine in the transverse direction after the subtraction of the local background extracted in the 10–20 Mpc region surrounding the filament. For illustration, the spectrum is binned to a minimum signal-to-noise ratio of 2.5. The left panel shows that the combination of a single temperature CIE (red dashed line), and the unmasked sources models (green, blue, and grey) can well fit most of the spectrum, except the significant excess at $E<0.6$ keV. The best-fit temperature of the WHIM gas is $0.58\pm 0.10$ keV. The right panel shows a fitting by a log-normal temperature distribution with $\log(\mu/\mathrm{K})=5.9$ and $\sigma=0.45$ dex. The 0.5–0.6 keV energy range can be well fitted by this model, though there is a likely deficit at $\sim 0.7$ keV.
@@ 5.2 Stacked spectrum and spectral fitting
We stacked the 0–10 Mpc transverse distance region spectra as the source spectrum and the 10–20 Mpc transverse distance region spectra as the local background region spectrum. The stacked background spectrum is smoothed using a Wiener filter with a kernel size of five channels for direct subtraction. Additionally, we compare the 6–9 keV count rate of the two spectra and store the ratio as the backscal in the header for further correction. The stacked net source spectrum is plotted in Fig. 9.
@@
We used XSPEC version 12.13.0 and the BXA [^5] https://github.com/JohannesBuchner/BXA package (Buchner et al. 2014) for our spectral analysis. The spectrum was first optimally binned (Kaastra & Bleeker 2016) using the tool ftgrouppha and then loaded into XSPEC. We used the 0.5–6.0 keV range for analysis and adopted C-stat (Kaastra 2017) to derive the fitting likelihood. The BXA package uses the nested sampling algorithm UltraNest (Buchner 2016, 2019, 2021) for Bayesian inference.
@@
We fit the spectrum using four components that account for the detected WHIM, unmasked halo gas, unmasked AGN, and unmasked XRB. The spectral model of the three unresolved sources is described in Sect. 4.1. We calculated the expected normalization ratios among the three components based on the modeled unresolved source fractions in Sect. 4.2 and coupled the ratios during fitting. To model the WHIM component, we used a single temperature (1-T) APEC model with fixed abundance $Z=0.2Z_{\sun}$. This value is based on the reported $\lesssim 0.3Z_{\odot}$ abundance in cluster outskirts by both observations and numerical simulations (Bulbul et al. 2016; Mernier et al. 2017; Biffi et al. 2017) and $>0.1Z_{\odot}$ in $z<1$ filaments from numerical simulations (Martizzi et al. 2019). This model combination can well fit the $E>0.7$ keV spectrum range with a best-fit parameter $kT=0.58\pm 0.10$ keV and a fitting statistics $c_{\mathrm{stat}}/\mathrm{d.o.f.}=105/111$. We tried to fix the metalicity to $0.1Z_{\sun}$ and $0.3Z_{\sun}$ to test the impact of the metalicity assumption and found there is no impact on the best-fit temperature. In addition, we also fitted a stacked 0–4 Mpc transverse distance region spectrum and obtained a similar $0.56\pm 0.09$ keV temperature.
@@
Recently, Vijayan & Li (2022) demonstrated that fitting a multi-temperature nature spectrum using a 1-T model will lead to an overestimation in the temperature. It explains that our best-fit temperature using a 1-T model does not agree with the majority of the gas phases around $10^{6}$ K predicted by simulations (e.g., Martizzi et al. 2019; Galárraga-Espinosa et al. 2021). In our stacked spectrum, we observe an excess feature in 0.5–0.6 keV, which agrees with the energy range of $\sim 0.56$ keV O VII lines emitted by cooler phases of gas around $10^{6.3}$ K ($\sim 0.17$ keV), suggesting the presence of cooler phases. The gas in these temperature phases only contributes to lines below 0.7 keV, while the fitting using a 1-T model tends to maximize the likelihood around the peak of the observed spectrum, which is in the 0.8–0.9 keV energy range, dominated by several Fe XVII lines emitted by $10^{6.8}$ K gas.
@@
We also tried to fit the spectrum using a multi-temperature model. We adopted a log-normal temperature distribution DEM model, with a peak at $\log(T/\mathrm{K})=5.9$ and $1\sigma$ width of 0.45 dex (as suggested by our analysis of numerical simulations; see more in Sect. 6.2). We used a simplified assumption of CIE condition and $0.2Z_{\sun}$ abundance of the WHIM. The log-normal DEM model can better fit the stacked spectrum in the 0.5–0.6 keV energy range (see the bottom panel of Fig. 9), though this fitting shows the feature of a likely deficit around 0.7 keV. This could be because the true temperature distribution is more complex than the log-normal distribution we use here, which creates discrepancies in some energy ranges.
6 Detected phases of WHIM
@@ 6.1 Surface brightness profile modeling
Table 3: Parameters and corresponding priors to constrain the density profile of the observed phase WHIM.
@@
With the best-fit temperature obtained in the previous section, we fit the density profiles of the gas in cosmic filaments. In this section, we use $r$ to denote the physical radius of gas cells or particles to the filament spine in the 3D space. Based on results from numerical simulations (Cui et al. 2019; Galárraga-Espinosa et al. 2021; Tuominen et al. 2021; Zhu et al. 2021), the gas in filaments has a multi-phase nature, ranging from $10^{3}$ to $10^{7}$ K in temperature and from $10^{-6}$ to $10^{-1}$ cm-3 in hydrogen density. Therefore, we introduced the parameter, $f_{\mathrm{vol}}$, to represent the fraction of the volume of the detected gas to the total filament volume. The volume-averaged emissivity of the detected phase gas at a certain radius is
@@
where $n_{\mathrm{H}}$ is the hydrogen density, 1.20 is the $n_{\mathrm{e}}/n_{\mathrm{H}}$ ratio for the plasma at the detected temperature. We used a King-type profile (King 1962), which is widely used to fit central flattened profiles, to model the radial change of both $n_{\mathrm{H}}$ and $f_{\mathrm{vol}}$,
@@
where $n_{\mathrm{H,0}}$ ($f_{\mathrm{vol,0}}$), $r_{\mathrm{c,n}}$ ($r_{\mathrm{c,fvol}}$), and $\alpha_{\mathrm{n}}$ ($\alpha_{\mathrm{fvol}}$) are the central normalization, core radius, and slope in the outskirts for $n_{\mathrm{H}}$ (or $f_{\mathrm{vol}}$), respectively. Because the radial dependencies of $n_{\mathrm{H}}$ and $f_{\mathrm{vol}}$ have a strong degeneracy, we assume that they have the same core radii, namely, $r_{\mathrm{c,n}}=r_{\mathrm{c,fvol}}$.
In addition, we must make several assumptions. First, we assume a constant baryon density contrast $\Delta_{\mathrm{b,0}}$ at the filament spine for filaments at all redshifts. In this case,
@@
where $\rho_{\mathrm{crit}}$ is the critical density at $z=0$, the factor 1.43 is the ratio between total mass density and hydrogen mass density, $m_{\mathrm{H}}$ is the hydrogen mass, and the redshift evolution comes from the expansion of the physical scale. Second, there is an outer boundary of the filaments $r_{\mathrm{out}}$, beyond which the gas density of our detected phase is zero. In this work, we have adopted a value of 10 Mpc. Third, the observed profile is smoothed and broadened due to the position uncertainty of the skeleton constructed from incomplete galaxy samples (e.g., Fig. 11 of Galárraga-Espinosa et al. 2021). Including these assumptions, the projected surface brightness at a given projected radial distance, $r_{\mathrm{proj}}$, is then
@@
where $l$ is depth along the line of sight; $K_{\mathrm{ph}}$ is the K-correction of photon spectrum given the spectral shape, redshift, and energy band; the term $(1+z)^{-3}$ is the photon dimming factor; $\Lambda$ is the cooling function calculated using the APEC model given the temperature, the assumption of $0.2Z_{\sun}$ abundance and the energy band; $A_{\mathrm{CF}}$ is the conversion factor from photon flux to count rate, which is calculated given the on-axis ARF, observation band, observed photon spectrum, and Galactic absorption curve.
When computing the fitting statistics, we constructed 7817 individual model surface brightness count rate profiles based on the redshift of each filament. We use the solid angle of each filament as the weight to obtain an averaged surface brightness profile as
@@
We used a Bayesian inference approach to obtain posterior distributions of the six parameters: $\log\Delta_{\mathrm{b,0}}$, $\log f_{\mathrm{vol,0}}$, $\log r_{\mathrm{c,n}}$, $\alpha_{\mathrm{n}}$, $\alpha_{\mathrm{fvol}}$, and $\log\sigma_{\mathrm{sm}}$, as listed in Table 3. In addition to the fitting statistics $\chi^{2}$, the likelihood function includes an additional constraint from SZ stacking results (Tanimura et al. 2020a), which is necessary to break up the degeneracy between $\Delta_{\mathrm{b,0}}$ and $f_{\mathrm{vol,0}}$. The predicted Compton-$y$ strength is
@@
where $\sigma_{\mathrm{T}}$ is the Thomson scattering cross section, $m_{\mathrm{e}}$ is the electron mass, and $c$ is the speed of light. The profiles of Tanimura et al. (2020a) show similar broadening effects as our stacked X-ray profile. Therefore, we also convolved the predicted Compton-$y$ profile with the same Gaussian kernel to create a smoothed Compton-$y$ profile as
@@
Given that the phases of the WHIM we detect only contribute a fraction of the total stacked SZ signal, we tentatively set the upper limit of the $\log y_{\mathrm{pd,conv}}(0)$ to be -7.5, which is approximately the stacked SZ signal at filament spine reported by (Tanimura et al. 2020a). We used a likelihood term $P(\log y_{\mathrm{pd,conv}}(0))=\mathcal{N}(\mu_{\mathrm{logy}},\sigma_{\mathrm{logy}})$ as the addition constraint. We do not know how much of the total stacked Compton-$y$ is contributed by the X-ray-emitting gas. Nevertheless, we adopt $\mu_{\mathrm{logy}}=-8.3$ and $\sigma_{\mathrm{logy}}=0.5$ in our analysis. The combined likelihood function is
@@
We used a Markov chain Monte Carlo (MCMC) method and the python package Zeus [^6] https://zeus-mcmc.readthedocs.io/en/latest/ (Karamanis et al. 2021) to perform Bayesian inference. The marginalized posterior distributions shows that the $\log\Delta_{\mathrm{b,0}}=2.28\pm 0.22$ and $\log f_{\mathrm{vol,0}}=-1.91\pm 0.28$. By using all MCMC chains for calculating the averaged baryon density contrast, we obtained $\log\bar{\Delta}_{\mathrm{b}}=1.88\pm 0.18$, which corresponds to a hydrogen density $4.4^{+2.2}_{-1.5}\times 10^{-5}$ cm-3 at the median redshift of the selected filament sample $z=0.47$. We note that we assumed a $Z=0.2Z_{\sun}$ metalicity for this inference. If the WHIM has a metalicity of $0.1Z_{\sun}$ (and $0.05Z_{\sun}$), at the temperature of 0.6 keV, the radiative cooling function in our analysis band is a factor of 1.5 (or 2.5) smaller than that of the $0.2Z_{\sun}$ assumption, which would introduce a 0.09 (and 0.2) dex increase in the inferred $\log\Delta_{\mathrm{b,0}}$.